the mathematical comparisons of energy/power loss of transmitting electricity at
different voltages. Produce a qualitative, reasoned argument which identifies why.
1
Expert's answer
2020-05-27T10:39:48-0400

Answer: The power loss in the line is calculated by the formula "P=I^2\times R" . The lower the current, the lower the power dissipation in the line. You can reduce the current and save power only by increasing the voltage, according to "P=U\times I."